ISRAEL JOURNAL OF MATHEMATICS

TABLE OF CONTENTS, VOLUME 242, 2021

Acciarri, C. and P. Shumyatsky A stronger form of Neumann’s BFC-theorem

269

Aistleitner, C. and S. Baker On the pair correlations of powers of real numbers

243

Akhtiamov, D., S. O. Ivanov and F. Pavutnitskiy Right exact localizations of groups

839

Baker, S. See C. Aistleitner Baldassarri, F. and V. Bojković Metric uniformization of

morphisms of Berkovich curves via p-adic differential equations

797

Barrera, D. and S. Gao Overconvergent Eichler–Shimura isomorphisms for unitary Shimura curves over totally real fields

707

Basu, R., S. Ganguly, M. Hegde and M. Krishnapur Lower deviations in β-ensembles and law of iterated logarithm in last passage percolation

291

Bojković, V. See F. Baldassarri Brian, W. R., A. Dow, D. Milovich and L. Yengulalp

Telgársky’s conjecture may fail

325

D’Aquino, P., A. Fornasiero and G. Terzo A weak version of the strong exponential closure

697

Dolphin, A. and A. Quéguiner-Mathieu The canonical quadratic pair on a Clifford algebra and triality

171

Dow, A. See W. R. Brian Entin, A. Factorization statistics of restricted polynomial specializa-

tions over large finite fields

37

Fornasiero, A. See P. D’Aquino Ganguly, S. See R. Basu Gao, S. See D. Barrera Goulart, V. and N. C. Saldanha Locally convex curves and the

Bruhat stratification of the spin group

565

Grimmelt, L. and W. Sawin Representation of squares by nonsingular cubic forms

501

Gruslys, V., S. Letzter and N. Morrison Lagrangians of hypergraphs II: When colex is best

637

Gügümcü, N. and S. Lambropoulou Braidoids  955

Hegde, M. See R. Basu Hrušák, M., C. A. Martínez-Ranero and U. A. Ramos-García

Hereditary interval algebras and cardinal characteristics of the continuum

769

Ivanov, S. O. See D. Akhtiamov Jain, V. Approximate Spielman–Teng theorems for the least singular

value of random combinatorial matrices

461

Krishnapur, M. See R. Basu Lambropoulou, S. See N. Gügümcü Lau, Y., M. H. Ng and Y. Wang On the exceptional set of the

generalized Ramanujan conjecture for GL(3)

279

Lee, J. and B. Schülke Convex graphon parameters and graph norms

549

Letzter, S. See V. Gruslys Levcovitz, I. Comparing the Roller and B(X) boundaries of CAT(0)

cube complexes

129

Ma, D. and K. Tasaka Relationships between multiple zeta values of depths 2 and 3 and period polynomials

359

Margolis, L. and O. Schnabel Corrigendum to “The twisted group ring isomorphism problem over fields”

997

Martínez-Ranero, C. A. See M. Hrušák Mikulincer, D. Stability of Talagrand’s Gaussian transport-entropy

inequality via the Föllmer process

215

Mildenberger, H. and S. Shelah The cofinality of the symmetric group and the cofinality of ultrapowers

97

Milovich, D. See W. R. Brian Minamoto, H. Homological identities and dualizing complexes of

commutative differential graded algebras

1

Molcho, S. Universal stacky semistable reduction  55
Moreno, J. P. Porosity and diametrical completeness  875

Morris, J. and P. Spiga Asymptotic enumeration of Cayley digraphs

401

Morrison, N. See V. Gruslys Ng, M. H. See Y. Lau

Park, C. A classification of Breuil modules  605

Pavutnitskiy, F. See D. Akhtiamov Quéguiner-Mathieu, A. See A. Dolphin Ramos-García, U. A. See M. Hrušák Saldanha, N. C. See V. Goulart Sawin, W. See L. Grimmelt Schnabel, O. See L. Margolis Schülke, B. See J. Lee Shelah, S. See H. Mildenberger Shi, R. Central limit theorem and cohomological equation on homo-

geneous spaces

891

Shumyatsky, P. See C. Acciarri Solecki, S. Transfinite sequences of topologies, descriptive

complexity, and approximating equivalence relations

933

Spiga, P. See J. Morris Tasaka, K. See D. Ma Terzo, G. See P. D’Aquino

Unger, S. Successive failures of approachability  663
Vávra, T. Periodic representations in Salem bases  83

Wang, Y. See Y. Lau Yengulalp, L. See W. R. Brian